[ The area of a circle is 27.5 square inches. To the nearest tenth of a square inch, find the area of the square that is circumscribed [around the outside] of this circle. ]
Found 2 solutions by TimothyLamb, rfer:
Answer by TimothyLamb(4379)

circle:
a = pi*r^2
r = sqrt( a/pi )
r = sqrt( 27.5/3.1415 )
r = 2.9586
d = 2r
d = 5.9173
---
square:
a = Lw
L = d
w = d
a = dd
a = 5.9173^2
a = 35.0141
---
Answer:
35.0 sq.in
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php

Answer by rfer(16322)

a=pir^2
27.5/3.14=8.758
sq rt 8.758=2.959
2.959*2=5.91877
----------
A=5.91877^2
A=35.0 sq in